OpenStudy (anonymous):

How is the integral of (2-2t)dt, equal to 2t-2^2 Thanks!

hartnn (hartnn):

using \(\huge \int t^n dt =\dfrac{t^{n+1}}{n+1}+c\)

hartnn (hartnn):

for 1st term , n=0 (2=2*t^0) for 2nd term., n=1

hartnn (hartnn):

bdw its 2t-t^2+c

OpenStudy (anonymous):

I have no idea whats going on, shouldn't this be a simple integral?

hartnn (hartnn):

it is. use of just 1 formula.

hartnn (hartnn):

using that formula, what will be integral of t dt ?

OpenStudy (anonymous):

Yea I know its the tried and true integral formula I must have done 100 of them, and I can't do this right now, my wires must be crossed.

OpenStudy (anonymous):

like the integral of x^2 is 1/3 x^3

OpenStudy (anonymous):

integral of t dt is 1/2 t^2

hartnn (hartnn):

yes, yes! correct.

hartnn (hartnn):

in same way, integral of 2 is 2t

OpenStudy (anonymous):

I guess I don't get that part

OpenStudy (anonymous):

like how do I just add the t in there

OpenStudy (anonymous):

I get that the derivative of 2t is 2

hartnn (hartnn):

not just adding 't' using the formula 2 = 2t^0 so, exponent = n= 0 after integration the exponent = n+1 = 0+1 =1 got this ?

OpenStudy (anonymous):

yes I do thank you so much because you probably just saved me an ulcer

hartnn (hartnn):

2 as a constant can be taken out of integral thats why 2t

OpenStudy (anonymous):

thats what I don't get

OpenStudy (anonymous):

if the 2 comes out, doesn't it remain a 2

OpenStudy (anonymous):

I'm thinking in derivative terms I guess where constants are dropped. with integral it goes out front so it never gets integrated?

OpenStudy (anonymous):

I guess its connected via subtraction to the 2t so it can't be pulled out

hartnn (hartnn):

integral of any constant = constant * variable integral of a dx = a* integral x^0 dx = ax+c

OpenStudy (anonymous):

ok thanks for your help! Im gonna let this marinate.

hartnn (hartnn):

welcome :)